# Hyperspherical Normalization for Scalable Deep Reinforcement Learning

### v1

@@ 1 Introduction
Over the past decade, a scaling law has emerged as a cornerstone in deep learning, suggesting that increasing the model size, compute, and data consistently enhance performance (Kaplan et al., 2020; Dehghani et al., 2023). This paradigm has driven significant breakthroughs in supervised learning (SL), including large language models (Gemini et al., 2023; Achiam et al., 2023) and diffusion models (Ramesh et al., 2021; Rombach et al., 2022).
In contrast, scaling laws often fail to apply to reinforcement learning (RL) (Song et al., 2019; Li et al., 2023). Unlike SL where data distributions remain static, RL agents are trained under evolving data distributions and shifting objectives (Sutton & Barto, 2018). This inherent non-stationarity impedes scaling efforts, as increasing the model capacity or computational resources causes the model to easily overfit to earlier data and reduces the agent’s adaptability to new tasks (Lyle et al., 2022; Dohare et al., 2023).
Figure 1: Compute vs RL Performance. Soft Actor-Critic with SimbaV2 outperforms other RL algorithms, where performance scales as compute increases. The grey numbers below each dot indicate the update-to-data (UTD) ratio. SimbaV2, with UTD = 1, achieves a performance of 0.848, surpassing TD-MPC2 (0.749), the most computationally intensive version of Simba (0.818), and BRO (0.807). The results show normalized returns, averaged over $57$ continuous control tasks from MuJoCo, DMC, MyoSuite, and HumanoidBench, each trained on 1 million samples.
@@
To address these challenges, the RL community has explored regularization techniques from SL. Methods such as weight decay (Farebrother et al., 2018) and dropout (Hiraoka et al., 2021) have been applied to reduce overfitting and improve generalization. Also, normalization techniques, such as layer normalization (Lei Ba et al., 2016; Lyle et al., 2023) and RL-specific variants (Bhatt et al., 2024; Lee et al., 2024c), have been employed to stabilize training by constraining the feature norm. More recently, periodic weight re-initialization (Nikishin et al., 2022; D’Oro et al., 2023; Schwarzer et al., 2023) has emerged as a strategy to address the instability in non-stationary optimization processes.
Despite their effectiveness, these approaches have notable limitations. Tuning regularizers is particularly challenging in RL’s dynamic settings; overly restrictive constraints impede training, while insufficient constraints fail to prevent overfitting (Henderson et al., 2018; Eimer et al., 2023). Moreover, normalization layers do not strictly regularize the growth of the parameter norm, resulting in varying effective learning rates (i.e., the gradient-to-parameter norm ratio) across layers and potentially destabilizing training (Lyle et al., 2024). Finally, periodic weight reinitialization becomes computationally prohibitive for large models or complex tasks (Sokar et al., 2023; Lee et al., 2024b).
@@
In response, we present SimbaV2, a novel RL architecture designed to stabilize non-stationary optimization by constraining weight, feature, and gradient norms. Building on the Simba architecture (Lee et al., 2024c), which uses pre-layernorm residual blocks (Xiong et al., 2020) and weight decay (Krogh & Hertz, 1991) to control the weight and feature norm growth, SimbaV2 introduces two modifications:
@@
  • Hyperspherical Normalization: We replace all layer normalization with hyperspherical normalization (i.e.,$\ell_{2}$-normalization) and project weights onto the unit-norm hypersphere after each gradient update. These changes ensure consistent effective learning rates across layers and eliminate the need for tuning weight regularization.
@@
  • Distributional Value Estimation with Reward Scaling: To address unstable gradient norms from varying reward scales and outliers, we integrate a distributional critic (Bellemare et al., 2017) and apply reward scaling to ensure the unit variance of target values for both actor and critic.
@@
Using soft actor-critic (Haarnoja et al., 2018) as a base, SimbaV2 stabilizes weight, feature, gradient norms, along with the effective learning rate throughout training (Section 5.2 and Figure 4). We evaluate SimbaV2 on four standard online RL benchmarks, MuJoCo (Todorov et al., 2012), DMC Suite (Tassa et al., 2018), MyoSuite (Caggiano et al., 2022), and HumanoidBench (Sferrazza et al., 2024), as well as D4RL MuJoCo benchmark (Fu et al., 2020) for offline RL. As indicated by results in Figures 1 and 2, SimbaV2 achieves state-of-the-art performance without algorithmic modifications or hyperparameter tuning. SimbaV2 also scales effectively with increased model size and computation, avoiding overfitting even in the absence of periodic reinitialization. We will release our code in the camera-ready version.
@@ 2.1 Regularization in Deep Reinforcement Learning
Deep RL is particularly susceptible to overfitting due to its inherently non-stationary optimization process (Song et al., 2019). To address overfitting, researchers have adapted regularization techniques from SL, including weight decay (Farebrother et al., 2018), dropout (Hiraoka et al., 2021), and various normalization layers (Gogianu et al., 2021; Bjorck et al., 2021; Lyle et al., 2023; Gallici et al., 2024; Bhatt et al., 2024; Lee et al., 2024c; Elsayed et al., 2024; Palenicek et al., 2025). However, these methods often prove insufficient when scaling RL models, as larger computational resources and increased model sizes can easily exacerbate overfitting (Li et al., 2023; Nauman et al., 2024a).
To further scale computations and model sizes in RL, recent studies have explored periodic weight reinitialization strategies in RL to rejuvenate learning and escape local minima (D’Oro et al., 2023; Nauman et al., 2024b). These strategies include reinitializing weights to their initial distributions (Nikishin et al., 2022), interpolating between random and current weights (Xu et al., 2023; Schwarzer et al., 2023), utilizing momentum networks (Lee et al., 2024b), and selectively reinitializing dormant weights (Sokar et al., 2023). While promising, reinitialization has a notable limitation: it can lead to the loss of useful information and incur significant computational overhead as model size increases.
@@ 3 Preliminaries
As background, we briefly explain Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018) and the Simba architecture (Lee et al., 2024c).
@@ 3.1 Soft Actor Critic
SAC is a prominent off-policy algorithm for continuous control. It aims to maximize both expected cumulative reward and policy entropy where $\tau=(o,a,r,o^{\prime})$ represents a transition tuple. SAC comprises a stochastic policy $\pi_{\theta}(a|o)$, a Q-function $Q_{\phi}(o,a)$, and an entropy coefficient $\alpha$ that balances reward and entropy.
The policy network is optimized to maximize the expected return while encouraging entropy which is formalized as:
@@
The Q-function $Q_{\phi}(o,a)$ is trained to minimize the Bellman residual:
@@ 5 Experiments
We now present a series of experiments designed to evaluate SimbaV2. Our investigation centers on three main setups:
  • Empirical Analysis (Section 5.2). Investigates whether SimbaV2 stabilizes the non-stationary optimization process and whether scaling model capacity and computation improves performance without overfitting.
@@
  • Design Study (Section 5.4.) Conducts ablation studies on individual architectural component of SimbaV2.
@@ 5.1 Experimental Setup
For the parameter scaling experiments, we focus on scaling the critic as prior studies suggest that scaling the actor has limited benefits (Nauman et al., 2024b; Lee et al., 2024c). Specifically, the critic’s hidden dimension is scaled from $\{128,256,512,1024\}$, growing is parameters from $0.3$M to $17.8$M. For computation scaling experiments, the UTD ratio is scaled from $\{1,2,4,8\}$ both with and without periodic weight reinitialization. We apply reinitialization every 500k update steps as suggested by Nauman et al. (2024b).
@@
5.2 Empirical Analysis
Training Dynamics. To understand the training dynamics of SimbaV2, we measure the feature norm, weight norm, gradient norm, and the effective learning rate (ELR), defined as the ratio of the gradient norm to the weight norm (Kodryan et al., 2022; Lyle et al., 2024) (See Appendix G for details). We weighted average each metric across layers where weights correspond to each layer’s fraction of total parameters. Additionally, we divide the layers into encoder layers (all layers before the output prediction) and predictor layers (those after) to analyze their respective dynamics.
Figure 5: Parameter Scaling. We scale the number of model parameters by increasing the width of the critic network. On DMC-Hard, both Simba and SimbaV2 benefit from increased model size. On HBench-Hard, however, Simba plateaus at larger model sizes, whereas SimbaV2 continues to improve.
Parameter Scaling. Next, we investigate whether SimbaV2’s stable training dynamics alleviate overfitting and enhance performance as the model size scales. In Figure 5, we increase the critic network’s parameters on DMC-Hard (left) and HBench-Hard (right). Both Simba and SimbaV2 gain from a larger model on DMC-Hard. However, on HBench-Hard, despite comparable performance at the smallest scale (0.3M), Simba plateaus at larger scales, whereas SimbaV2 continues to improve. This suggests that stabilizing training dynamics can effectively leverage a larger model.
Compute Scaling. Finally, we explore compute scaling, a major factor in improving sample efficiency in deep RL (Li et al., 2023). While increasing the UTD ratio can enhance sample efficiency, it also raises the risk of overfitting. Previous work has addressed this with ensembling (Chen et al., 2021b), periodic reinitialization (Lee et al., 2024a; D’Oro et al., 2023; Nauman et al., 2024b), or both (Kim et al., 2023). We investigate whether SimbaV2’s stable dynamics enable effective scaling without these additional methods.
Figure 6: Compute Scaling.. We scale compute by increasing the UTD ratio. We compare Simba and SimbaV2, both with and without periodic reset. Simba saturates at lower ratios without reset but improves with reset. In contrast, SimbaV2 scales smoothly even without reset, where using reset slightly degrades its performance.
@@ 5.2 Empirical Analysis
To assess the importance of hyperspherical normalization for UTD scaling, we test a variant in Appendix H that adds distributional critics and reward scaling but omits the normalization. This variant fails to scale at higher UTD ratios, demonstrating the role of hyperspherical normalization.
@@ Appendix C Hyperparameters
Number of atoms $n_{atoms}$ 101
@@ Appendix E Baselines
SAC (Haarnoja et al., 2018). Soft Actor-Critic (SAC) is an off-policy actor-critic algorithm in which the actor simultaneously maximizes expected return and entropy, encouraging both stability and exploration. Results for Gym-MuJoCo were obtained from Fujimoto et al. (2023). These scores are average over $10$ random seeds. For DMC, MyoSuite, and HBench tasks, we use the results from Lee et al. (2024c) which were obtained by running the official repository for 10 random seeds.
@@
CrossQ (Bhatt et al., 2024). CrossQ achieves superior performance and sample efficiency with low replay ratio, by removing target networks and employing careful batch normalization. Results for Gym-MuJoCo were obtained by running the official repository (https://github.com/adityab/CrossQ) for 10 random seeds.
@@
Simba (Lee et al., 2024c). SimBa is an architecture designed to scale up parameters in deep reinforcement learning by injecting a simplicity bias with observation normalizer, residual blocks, and layer normalizations. For Gym-MuJoCo, DMC, MyoSuite, and HBench tasks, we use the results from Lee et al. (2024c) which were obtained by running the official repository (https://github.com/SonyResearch/simba) over $15$ random seeds for DMC hard tasks and $10$ random seeds otherwise.

### v2

@@ 1 Introduction
Over the past decade, a scaling law has emerged as the cornerstone of supervised learning (SL), suggesting that increasing model size, compute, and data consistently improve performance (Kaplan et al., 2020; Dehghani et al., 2023). This paradigm has driven significant breakthroughs, from large language models (Gemini et al., 2023; Achiam et al., 2023) to diffusion models (Ramesh et al., 2021; Rombach et al., 2022), where bigger models reliably translate to better performance.
In contrast, scaling laws often fail to apply in reinforcement learning (RL) (Song et al., 2019; Li et al., 2023). Unlike SL’s static data distributions, RL agents must contend with continuously evolving data distributions and shifting objectives throughout their training process (Sutton & Barto, 2018). This fundamental non-stationarity creates a scaling paradox: increasing model capacity or computational resources frequently leads to overfitting to earlier experiences and reduced adaptability to new tasks (Lyle et al., 2022; Dohare et al., 2023).
Figure 1: Compute vs RL Performance. Performance scales with increased compute when using Soft Actor-Critic with SimbaV2 architecture, outperforming other state-of-the-art RL algorithms. SimbaV2 achieves 0.848 normalized return with an update-to-data (UTD) ratio of 1, surpassing TD-MPC2 (0.749 at UTD=1), Simba (0.818 at UTD=8), and BRO (0.807 at UTD=8). Grey numbers below each point indicate the UTD ratio. Results are averaged over 57 continuous control tasks from MuJoCo, DMC, MyoSuite, and HumanoidBench, each trained on 1 million samples.
@@
One of the root causes of RL’s scaling challenges lies in uncontrolled norm growth during training, which destabilizes optimization in multiple ways. Feature norms grow uncontrollably due to the implicit bias of TD loss (Kumar et al., 2022), where dominant dimensions cause overfitting and loss of plasticity (Lyle et al., 2022; Ma et al., 2023). Parameter norms grow unbounded, reducing effective learning rates (gradient-to-parameter ratio) and making weight updates increasingly difficult (Dohare et al., 2023; Lyle et al., 2024). Gradient norms fluctuate due to varying reward scales and outliers, further disrupting optimization. These instabilities are compounded with an increased model size or update frequency, making RL harder to scale than SL.
Previous work has addressed these norm instabilities through separate, isolated approaches. Normalization layers such as $\ell_{2}$-normalization (Bjorck et al., 2021; Hussing et al., 2024), layer normalization (Lei Ba et al., 2016; Lyle et al., 2023), and RL-specific variants (Bhatt et al., 2024; Lee et al., 2024c) control the growth of feature norm. Weight decay (Farebrother et al., 2018) manages the growth of parameter norm. Reward scaling and cross-entropy loss (Schaul et al., 2021; Farebrother et al., 2024) were adopted to control gradient norm fluctuations. However, these techniques are applied individually without a unified framework, making coordination and scaling difficult. Periodic weight reinitialization (Nikishin et al., 2022; D’Oro et al., 2023; Schwarzer et al., 2023) offers an alternative by completely retraining networks periodically. While effective, this approach requires additional training time and causes sharp performance drops, making it impractical for safety-critical applications.
@@
In response, we present SimbaV2, a novel RL architecture that addresses these challenges by simultaneously stabilizing weight, feature, and gradient norms within a unified framework. Building on the Simba architecture (Lee et al., 2024c), which uses pre-layernorm residual blocks (Xiong et al., 2020) and weight decay (Krogh & Hertz, 1991), SimbaV2 introduces three key modifications:
  • Hyperspherical Feature Normalization: We replace all layer normalization with hyperspherical normalization ($\ell_{2}$-normalization).
@@
  • Hyperspherical Weight Normalization: We remove weight decay and instead project weights onto the unit-norm hypersphere after each gradient update (Loshchilov et al., 2024). Combined with hyperspherical feature normalization, this ensures consistent effective learning rates across layers and eliminates the need for weight regularization tuning.
@@
  • Distributional Value Estimation with Reward Scaling: To address unstable gradient norms caused by varying reward scales and outliers, we integrate a distributional critic (Bellemare et al., 2017) and apply reward scaling to maintain unit variance of the target throughout training.
@@
Using Soft Actor-Critic (Haarnoja et al., 2018) as our base algorithm, SimbaV2 effectively stabilizes all three types of norms while maintaining consistent effective learning rates throughout training (Section 5.2 and Figure 4). We evaluated SimbaV2 on four standard online RL benchmarks: MuJoCo (Todorov et al., 2012), DMC Suite (Tassa et al., 2018), MyoSuite (Caggiano et al., 2022), and HumanoidBench (Sferrazza et al., 2024); as well as the D4RL MuJoCo benchmark (Fu et al., 2020) for offline RL. As shown in Figures 1 and 2, SimbaV2 achieves state-of-the-art performance without requiring algorithmic modifications or hyperparameter tuning, and scales effectively with increased model size and computation without using periodic reinitialization.
@@ 2.1 Regularization in Deep Reinforcement Learning
Deep RL is particularly susceptible to overfitting due to its inherently non-stationary optimization process (Song et al., 2019). To address overfitting, researchers have adapted regularization techniques from SL, including weight decay (Farebrother et al., 2018), dropout (Hiraoka et al., 2021), various normalization layers (Gogianu et al., 2021; Bjorck et al., 2021; Lyle et al., 2023; Gallici et al., 2024; Bhatt et al., 2024; Lee et al., 2024c; Elsayed et al., 2024; Palenicek et al., 2025), and mixture of expert (Obando-Ceron et al., 2024; Willi et al., 2024). However, these methods often prove insufficient when scaling RL models, as larger computational resources and increased model sizes can easily exacerbate overfitting (Li et al., 2023; Nauman et al., 2024a).
To further scale computations and model sizes in RL, recent studies have explored periodic weight reinitialization strategies to rejuvenate learning and escape local minima (D’Oro et al., 2023; Nauman et al., 2024b). These strategies include reinitializing weights to their initial distributions (Nikishin et al., 2022), interpolating between random and current weights (Xu et al., 2023; Schwarzer et al., 2023), utilizing momentum networks (Lee et al., 2024b), and selectively reinitializing dormant weights (Sokar et al., 2023). While promising, reinitialization has a notable limitation: it can lead to the loss of useful information and incur significant computational overhead as model size increases.
@@ 3 Preliminaries
As background, we briefly explain the Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018) and the Simba architecture (Lee et al., 2024c).
@@ 3.1 Soft Actor Critic
SAC is a prominent off-policy algorithm for continuous control. It aims to maximize both expected cumulative reward and policy entropy, where $\tau=(o,a,r,o^{\prime})$ represents a transition tuple. SAC comprises a stochastic policy $\pi_{\theta}(a|o)$, a Q-function $Q_{\phi}(o,a)$, and an entropy coefficient $\alpha$ that balances reward and entropy.
The policy network is optimized to maximize the expected return while encouraging entropy, which is formalized as:
@@
The Q-function $Q_{\phi}(o,a)$ is trained to minimize the Bellman residual loss:
@@ 4 SimbaV2
  • Weight Decay $\bm{\rightarrow}$ Weight Projection: Weight decay is replaced with direct weight projection onto the unit hypersphere after each gradient update.
@@ 5 Experiments
We now present a series of experiments designed to evaluate SimbaV2. Our investigation centers on four main setups:
  • Optinmization Analysis (Section 5.2). Investigate whether SimbaV2 stabilizes the optimization process.
  • Scaling Analysis (Section 5.3). Investigate whether SimbaV2 allows scaling model capacity and computation.
@@
  • Design Study (Section 5.5.) Conducts ablation studies on individual architectural components of SimbaV2.
@@ 5.1 Experimental Setup
5.2 Optimization Analysis
To understand the optimizatiom dynamics of SimbaV2, we measure the feature norm, weight norm, gradient norm, and the effective learning rate (ELR), defined as the ratio of the gradient norm to the weight norm (Kodryan et al., 2022; Lyle et al., 2024) (See Appendix G for details). We weighted average each metric across layers where weights correspond to each layer’s fraction of total parameters. Additionally, we divide the layers into encoder layers (all layers before the output prediction) and predictor layers (those after) to analyze their respective dynamics.
Figure 5: Width Scaling. We scale the number of model parameters by increasing the width of the critic network. On DMC-Hard, both Simba and SimbaV2 benefit from increased model size. On HBench-Hard, however, Simba plateaus at larger model sizes, whereas SimbaV2 continues to improve.
5.3 Scaling Analysis
For this experiment, we investigate whether SimbaV2’s stable training dynamics enable better scaling performance as model parameters or computational resources increase, while reducing overfitting compared to existing methods.
Experimental Setup. We conduct two types of scaling experiments. For parameter scaling, we focus on scaling the critic network, as prior studies indicate that scaling the actor provides limited benefits (Nauman et al., 2024b; Lee et al., 2024c). We test two scaling approaches: width scaling by varying the critic’s hidden dimension across $\{128,256,512,1024\}$, increasing parameters from $0.3M$ to $17.8M$; and depth scaling by varying the number of critic blocks $L$ across $\{1,2,4,8\}$, growing parameters from $2.2M$ to $17.8M$.
For compute scaling experiments, we vary the update-to-data (UTD) ratio across $\{1,2,4,8\}$. We compare results both with and without periodic weight reinitialization, since prior work suggests that compute scaling requires periodic reinitialization to avoid overfitting (D’Oro et al., 2022). Following Nauman et al. (2024b), we apply reinitialization every 500,000 update steps when used.
Parameter Scaling. Figure 5 shows width scaling results on DMC-Hard (left) and HBench-Hard (right). Both Simba and SimbaV2 benefit from larger models on DMC-Hard. However, on the more challenging HBench-Hard benchmark, while both methods achieve comparable performance at the smallest scale ($d_{h}=128$), their scaling behavior diverges significantly. Simba plateaus at larger scales with peak performance at $d_{h}=1024$, while SimbaV2 continues to improve with increased width. This demonstrates that SimbaV2’s stabilized training dynamics effectively leverages larger model capacity.
Figure 6 presents depth scaling results. In HBench-Hard, SimbaV2 shows consistent performance improvements as the depth of the critic $L$ increases, successfully solving the five complex tasks in $L=8$. On DMC-Hard, SimbaV2’s performance also improves with depth but begins saturating around $L=4$, likely due to task complexity limitations rather than architectural constraints. In contrast, Simba’s performance either plateaus around $L=2$ or slightly decreases after initial improvement. This clear difference demonstrates SimbaV2’s superior depth scalability, which we attribute to its effective regularization mechanisms that enable stable training of deeper networks.
Compute Scaling. We next explore compute scaling through increased UTD ratios, a key factor in improving sample efficiency in deep RL (Li et al., 2023). While higher UTD ratios can enhance sample efficiency, they also increase the risk of overfitting. Previous approaches address this through ensembling (Chen et al., 2021b), periodic reinitialization (Lee et al., 2024a; D’Oro et al., 2023; Nauman et al., 2024b), or both (Kim et al., 2023). We investigate whether SimbaV2’s stable dynamics enable effective scaling without these additional mechanisms.
Figure 6: Depth Scaling. We scale the number of model parameters by increasing the depth of the critic network. On both DMC-Hard and HBench-Hard, SimbaV2 benefits from increased depth, while Simba’s performance degrades beyond a shallow configuration of $L\geq 2$.
@@ 5.3 Scaling Analysis
To verify the importance of hyperspherical weight and feature normalization for UTD scaling, we test a variant in Appendix H.1 that includes distributional critics and reward scaling on Simba. This variant fails to scale at higher UTD ratios, confirming the critical role of hyperspherical normalization for effective scaling.
Figure 7: Compute Scaling.. We scale compute by increasing the UTD ratio. We compare Simba and SimbaV2, both with and without periodic reset. Simba saturates at lower ratios without reset, but improves with reset. In contrast, SimbaV2 scales smoothly even without reset, where using reset slightly degrades its performance.
@@ 6 Lessons and Opportunities
Impact Statement
This paper presents work aimed at advancing the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Acknowledgements
This work was supported by Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST)). This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2025-00555621) This work was mainly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2021-II212068, Artificial Intelligence Innovation Hub).
@@ Appendix C Hyperparameters
Number of atoms $n_{\text{atoms}}$ 101
@@ Appendix E Baselines
SAC (Haarnoja et al., 2018). Soft Actor-Critic (SAC) is an off-policy actor-critic algorithm in which the actor simultaneously maximizes expected return and entropy, encouraging both stability and exploration. For the MuJoCo tasks, results averaged over 10 random seeds were obtained directly from the Bhatt et al. (2024) authors, with the update-to-data (UTD) ratio set to 1. For DMC, MyoSuite, and HBench tasks, we use the results from Lee et al. (2024c) which were obtained by running the official repository for 10 random seeds, with the update-to-data (UTD) ratio set to 2.
@@
REDQ (Chen et al., 2021b). Randomized Ensembled Double Q-Learning (REDQ) expands clipped double Q-learning from two Q-networks to an ensemble of ten to control estimation bias and variance, and enhance training stability. For the MuJoCo tasks, results averaged over 10 random seeds were obtained directly from the Bhatt et al. (2024) authors, with the update-to-data (UTD) ratio set to 20.
DroQ (Chen et al., 2021b). Dropout Q-Function (DroQ) reduces the computational burden of REDQ by using a smaller ensemble of Q functions while employing Dropout and Layer Normalization to stabilize training against Dropout-induced noise. For the MuJoCo tasks, results averaged over 10 random seeds were obtained directly from the Bhatt et al. (2024) authors, with the update-to-data (UTD) ratio set to 20.
@@
CrossQ (Bhatt et al., 2024). CrossQ achieves superior performance and sample efficiency with low replay ratio, by removing target networks and employing careful batch normalization. Results for Gym-MuJoCo were obtained by running the official repository (https://github.com/adityab/CrossQ) for 10 random seeds,
@@
Simba (Lee et al., 2024c). SimBa is an architecture designed to scale up parameters in deep reinforcement learning by injecting a simplicity bias with observation normalizer, residual blocks, and layer normalizations. For Gym-MuJoCo, DMC, MyoSuite, and HBench tasks, we use the results from Lee et al. (2024c) which were obtained by running the official repository (https://github.com/SonyResearch/simba) over $15$ random seeds for DMC hard tasks and $10$ random seeds otherwise. Unless stated otherwise, we set update-to-data (UTD) ratio to be $2$.